A ball is rolled on a flat surface with an initial velocity of 2 m/s. what is the ball's acceleration if it was found 15 meters away from its starting position after 5 seconds? What is the ball's velocity at a distance 20m? how long will the ball travel to achieve a velocity of 20 m/s?

1
Expert's answer
2022-02-23T08:31:50-0500

(a)

"d=v_0t+at^2\/2"

"a=2(d-v_0t)/t^2=2(15-2*5)/5^2=0.4\:\rm m/s^2"

(b)

"v=v_0+at"

"t=\frac{v-v_0}{a}=\frac{20-2}{5}=3.6\:\rm s"